### v1

@@ Abstract
Current benchmarks for AI clinician systems, often based on multiple-choice exams or manual rubrics, fail to capture the depth, robustness, and safety required for real-world clinical practice. To address this, we introduce the GAPS framework, a multidimensional paradigm for evaluating Grounding (cognitive depth), Adequacy (answer completeness), Perturbation (robustness), and Safety. Critically, we developed a fully automated, guideline-anchored pipeline to construct a GAPS-aligned benchmark end-to-end, overcoming the scalability and subjectivity limitations of prior work. Our pipeline assembles an evidence neighborhood, creates dual graph and tree representations, and automatically generates questions across G-levels. Rubrics are synthesized by a DeepResearch agent that mimics GRADE-consistent, PICO-driven evidence review in a ReAct loop. Scoring is performed by an ensemble of large language model (LLM) judges. Validation confirmed our automated questions are high-quality and align with clinician judgment. Evaluating state-of-the-art models on the benchmark revealed key failure modes: performance degrades sharply with increased reasoning depth (G-axis), models struggle with answer completeness (A-axis), and they are highly vulnerable to adversarial perturbations (P-axis) as well as certain safety issues (S-axis). This automated, clinically-grounded approach provides a reproducible and scalable method for rigorously evaluating AI clinician systems and guiding their development toward safer, more reliable clinical practice.
@@ 2.1 Evidence-Anchored Benchmark Construction
We implemented a multi-agent pipeline that converts guideline-centered evidence corpora into evaluated GAPS items with stable, automatically generated rubrics. The approach is guideline-agnostic: any clinical practice guideline can serve as an anchor, expanded through a bounded evidence neighborhood obtained by traversing backward citations up to $k$ hops. In this study, we instantiate the framework using one representative guideline with $k=3$ to illustrate the process.
@@
This modular, agentic pipeline generalizes across clinical domains and guidelines without requiring methodological adjustments. Using this framework, we generated a large corpus of automatically evaluated GAPS items. In this study, we present a curated subset of 92 questions,GAPS-NCCN-NSCLC-preview, constructed from the evidence neighborhood of the NCCN Non–Small Cell Lung Cancer (NSCLC) guideline. This subset exemplifies the framework’s scalability, auditability, and clinical grounding.
@@ 2.2 GAPS Evaluation Benchmark and Model Performance
We next evaluate representative state-of-the-art LLMs on the automatically constructed GAPS benchmark GAPS-NCCN-NSCLC-preview using rule-based and rubric-based scoring (normalized to [0,1]), details of which are covered in Section 3.6. Scores were obtained by an ensemble of LLM-as-a-judge models, with clinician audits on stratified subsets for validation. Performance was analyzed across the four GAPS axes: Grounding levels (G1–G4), Perturbations (P1–P3), and Adequacy and Safety outcomes (hit rates). The evaluated models included GPT-5 [12], Gemini 2.5 Pro [13], Claude Opus 4 [14], DeepSeek-V3.1-Terminus [15], and Qwen3-235B-A22B-Instruct-2507 [16].
@@ 3.6.2 LLM-as-a-Judge Ensemble
Element-wise scoring is performed by an ensemble of large language models acting as impartial judges, following frameworks such as HealthBench [8]. This setup blends rubric-based reasoning for A/S elements with rule-based checks for S1 relevance.

### v2

@@ Abstract
Current benchmarks for AI clinician systems, often based on multiple-choice exams or manual rubrics, fail to capture the depth, robustness, and safety required for real-world clinical practice. To address this, we introduce the GAPS framework, a multidimensional paradigm for evaluating Grounding (cognitive depth), Adequacy (answer completeness), Perturbation (robustness), and Safety. Critically, we developed a fully automated, guideline-anchored pipeline to construct a GAPS-aligned benchmark end-to-end, overcoming the scalability and subjectivity limitations of prior work. Our pipeline assembles an evidence neighborhood, creates dual graph and tree representations, and automatically generates questions across G-levels. Rubrics are synthesized by a DeepResearch agent that mimics GRADE-consistent, PICO-driven evidence review in a ReAct loop. Scoring is performed by an ensemble of large language model (LLM) judges. Validation confirmed our automated questions are high-quality and align with clinician judgment (90% agreement, Cohen’s Kappa 0.77). Evaluating state-of-the-art models on the benchmark revealed key failure modes: performance degrades sharply with increased reasoning depth (G-axis), models struggle with answer completeness (A-axis), and they are highly vulnerable to adversarial perturbations (P-axis) as well as certain safety issues (S-axis). This automated, clinically-grounded approach provides a reproducible and scalable method for rigorously evaluating AI clinician systems and guiding their development toward safer, more reliable clinical practice. The benchmark dataset GAPS-NSCLC-preview and evaluation code are publicly available at https://huggingface.co/datasets/AQ-MedAI/GAPS-NSCLC-preview and https://github.com/AQ-MedAI/MedicalAiBenchEval.
@@ 2 Results
Figure 1: Overview of the automated GAPS benchmark construction pipeline. The workflow operates in four stages: (1) Data Structuring: assembling a frozen evidence neighborhood to construct dual Knowledge Graph (KG) and Hierarchical Tree representations; (2) Topic Tree Construction: mapping the coverage space for systematic navigation; (3) Item Generation: synthesizing questions across Grounding levels (G1–G4) and generating robust variants via the Perturbation module (P0–P3); and (4) Rubric Synthesis: employing a GRADE-aware DeepResearch agent to formulate structured Adequacy (A) and Safety (S) rubrics.
@@ 2.1 Evidence-Anchored Benchmark Construction
We implemented a multi-agent pipeline (Figure 1) that converts guideline-centered evidence corpora into evaluated GAPS items with stable, automatically generated rubrics.
@@
This modular, agentic pipeline generalizes across clinical domains and guidelines without requiring methodological adjustments. Using this framework, we generated a large corpus of automatically evaluated GAPS items. In this study, we present a curated subset of 92 questions and 1,691 individual rubrics, GAPS-NSCLC-preview, constructed from the evidence neighborhood of the NCCN Non–Small Cell Lung Cancer (NSCLC) guideline. This subset exemplifies the framework’s scalability, auditability, and clinical grounding.
@@ 2.2 GAPS Evaluation Benchmark and Model Performance
We next evaluate representative state-of-the-art LLMs on the automatically constructed GAPS benchmark GAPS-NSCLC-preview using rule-based and rubric-based scoring (normalized to [0,1]), details of which are covered in Section 3.6. Scores were obtained by an ensemble of LLM-as-a-judge models, with clinician audits on stratified subsets for validation. Performance was analyzed across the four GAPS axes: Grounding levels (G1–G4), Perturbations (P1–P3), and Adequacy and Safety outcomes (hit rates). The evaluated models included GPT-5 [12], Gemini 2.5 Pro [13], Claude Opus 4 [14], DeepSeek-V3.1-Terminus [15], and Qwen3-235B-A22B-Instruct-2507 [16].
@@ 3.6 Rubric-Based Scoring
Figure 6: The automated clinical answer evaluation and analysis pipeline. The system processes response data through five layers: (1) Data Processing: parsing evaluation units into atomic claims; (2) Judge Model Invocation: employing an ensemble of parallel LLMs (e.g., GPT-4, Claude, DeepSeek) to independently evaluate Positive (Adequacy) and Negative (Safety) claims using deterministic sampling; (3) Aggregation: fusing judgments via majority voting and consistency checks; (4) Scoring: applying the hybrid rule-based logic with strict safety overrides for critical errors; and (5) Visualization: generating performance analytics across the GAPS axes.
@@ 3.6.2 LLM-as-a-Judge Ensemble
Element-wise scoring is performed by an ensemble of large language models acting as impartial judges (Layer 2 in Figure 6), following frameworks such as HealthBench [8]. This setup blends rubric-based reasoning for A/S elements with rule-based checks for S1 relevance.
@@
3.6.3 Reliability Validation
To ensure the trustworthiness of the automated evaluation, we rigorously validated the framework against independent annotations from five senior clinical experts. Across the GAPS benchmark GAPS-NSCLC-preview of 92 questions and 1,691 individual rubrics, the automated pipeline achieved an overall agreement rate of 90.00% with expert consensus. The Cohen’s Kappa coefficient [21] was 0.77, indicating “substantial agreement”. These results confirm that the GAPS automated judge possesses expert-level reliability.
@@ 5 Conclusion
Data Availability
The benchmark dataset constructed for this study (GAPS-NSCLC-preview) is available for research purposes at https://huggingface.co/datasets/AQ-MedAI/GAPS-NSCLC-preview.
Code Availability
The source code for the evaluation scripts is publicly available on GitHub at https://github.com/AQ-MedAI/MedicalAiBenchEval. The automated benchmark construction pipeline and the DeepResearch rubric agent will be made available at a later date.